OpenStudy (anonymous):

An electron moves around a proton in a circular orbit of radius 5.29*10^-11m. If the prton is m=1.67*10^-27kg. and the electron mass is m=9.11*10^-31kg.calculate the eletric and gravitational force exerted on the electron by the proton

OpenStudy (anonymous):

\[F_E = \frac{1}{4\pi\epsilon_0}\frac{q_1q_2}{r^2}\] \[q_e = e\] \[q_p = +e\] \[F_G = G\frac{m_pm_e}{r^2}\]

OpenStudy (anonymous):

^_^